# Preconditioned Score-based Generative Models

### v2

@@ 7.2 Evaluation on running speed
We evaluate the effect of running speed from using PDS. In this test, we use one NVIDIA RTX 3090 GPU. We track the average wall-clock time of generating a batch of 8 images. As shown in Table V, PDS only introduces marginal extra time complexity per iteration, corroborating our complexity analysis in Sec. 5.2.3. Importantly, PDS enables low-iteration generation for better efficiency. For example, for the high-resolution generation task on FFHQ ($1024$), PDS reduces the time cost of generating $8$ human facial samples from $1920$ to $68$ in seconds, i.e., 28$\times$ speedup.
TABLE V: Comparing the running time per iteration of generating a batch of $8$ images. Time unit: Seconds.
@@ 7.3 Relation between $\alpha$ and $T$
We test CIFAR-10 with only the frequency preconditioner and CelebA, ImageNet, and COCO ($64\times 64$) with both. The results of linear regression are depicted in Fig. 5. We obtain the coefficient of determination of 0.895 on CIFAR-10, 0.940 on CelebA($64\times 64$), 0.992 on ImageNet ($64\times 64$), and 0.804 respectively. This suggests strong linear relations, validating our analysis. Note, when $(1+\frac{1}{\alpha})^{2}-1$ and $(1+\frac{1}{\alpha})-1$ are similar, the ratio of the slope $a$ will be close to the square root of the dimension ratio. For instance, the relation between CIFAR-10 and CelebA($64\times 64$) fits as:
$\frac{a_{1}}{a_{2}}=\frac{28.53}{50.67}=0.56\sim 0.5=\sqrt{\frac{3\times 32\times 32}{3\times 64\times 64}}=\sqrt{\frac{d_{1}}{d_{2}}},$ (51)
@@
Figure 5: The relation between the iterations $T$ and the parameter $\alpha$. both suggesting strong linear correlation.
@@ 7.4 Sensitivity of parameters
We evaluate the parameter sensitivity of PDS. We perturb the optimal $\alpha$ (Table I) by ±$\pm 10\%$. As shown in Table VI, the performance is not parameter-sensitive.
TABLE VI: Parameter ($\alpha$) sensitivity analysis on CIFAR-10.
@@
7.5 Evaluating the skew-symmetric operator
We study the effect of the solenoidal term $S\bigtriangledown_{\mathbf{x}}\log p_{\mathbf{x}^{\ast}}(\mathbf{x})dt$ to the Langevin dynamics. As proved in Thm. 2, as long as $S$ is skew-symmetric, it will not change the steady-state distribution of the original process. To verify this, we generalize the original process as
@@ 7.5 Evaluating the skew-symmetric operator
where $\omega$ controls the weight of $S$. We experiment with $S[\cdot]=\mathrm{Re}[F[\cdot]-F^{\mathrm{T}}[\cdot]]$ for being skew-symmetric and $\omega\in[1,1000]$. Fig. 6 shows that $\omega$ has no impact to the generation quality. This verifies that $S$ does not change the steady-state distribution of the original Langevin dynamics.
———-$\omega=1$——–$\omega=10$——–$\omega=100$——–$\omega=1000$
Figure 6: Samples produced by NCSN++ [4] w/ PDS on FFHQ ($1024$x1024102410241024) with different solenoidal terms. Iterations: $66$.
We further perform similar tests with different iterations and other different skew-symmetric operator $S$:
@@
where $P_{m,n}$ is the shift operator that rolls the input image for $m$ places along the height coordinate and rolls the input image for $n$ places along the width coordinate. The sampling results are shown in Fig. 9, where we set $\omega=1000$. It is observed that again all these solenoidal terms do not impose an obvious effect on the sampling quality. Additionally, as displayed in Fig. 15 in the supplementary material, these solenoidal terms also do not make an obvious effect on acceleration. We leave more extensive study of the solenoidal terms in the further work.
7.6 Experiments on other SGMs
We evaluate the performance of PDS on two more SGMs, i.e., NCSN [2] and NCSNv2 [3]. Although the approach of constructing preconditioning operators in Sec. 5.2 works well for accelerating NCSN++ [4], it is less effective for NCSN [2] and NCSNv2 [3]. A possible reason is that the two models are less capable of exploiting the data statistics. To address this, we construct a simpler version of frequency preconditioner as
$R_{f}(c,h,w)=\left\{\begin{matrix}1,\mathrm{if}\quad(h-0.5H)^{2}+(w-0.5W)^{2}\leq 2r^{2}\\ \lambda,\mathrm{if}\quad(h-0.5H)^{2}+(w-0.5W)^{2}>2r^{2}\end{matrix}\right.,$ (60)
$M_{f}[\mathbf{x}]=F^{-1}[F[\mathbf{x}]\bullet R_{f}].$ (61)
Note that for NCSN and NCSNv2 have only corrector in the sampling process. We set $\lambda$ so that the operator $M_{f}$ balances the coordinates of low and high frequency components of its input via shrinking the scale of low-frequency coordinates. This is because that the scale of low-frequency coordinates is higher than that of high-frequency coordinates among natural images [54]. As shown in Fig. 1-3 in Appendix Fig. 4, PDS can dramatically accelerate the sampling process of NCSN on MNIST that of NCSNv2 on LSUN.
7.7 Compared with subspace diffusion generative models (SDGMs)
The subspace diffusion generative models (SDGMs) [45] restrict the sampling process of the SGMs on a low-dimensional subspace for acceleration. This might also help to mitigate the issue of ill-conditioning. We thus compare our PDS with SDGMs in terms of the effect of alleviating ill-conditioning.
We conduct two separate experiments to evaluate their performance. In the first experiment, we investigate the performance of two methods in terms of varying iterations. As shown in Table VII, our PDS achieves lower/better FID scores when compared to SDGMs subject to the same iterations. It is important to highlight that SDGM benefits from a smaller per-iteration cost due to reduced dimensionality, allowing it to execute more iterations within the same runtime. Hence we conduct a second experiment with a comparison taking into account the runtime and sample quality of the two methods, aiming to maintain a different fair evaluation. As depicted in Fig. 7, our PDS demonstrates the ability to generate higher-quality images (as indicated by a lower FID score) at the equivalent runtime cost when compared to SDGMs. In summary, our preconditioning strategy surpasses the dimension reduction idea of SDGM.
Iterations $T$ 50 100 200 400 500 1000
PDS 4.90 2.90 2.53 2.19 2.11 1.99
SDGM 386.70 64.05 7.03 3.03 2.70 2.17
TABLE VII: Comparison of our PDS and SDGM [45] on CIFAR-10 generation under varying iterations.
Figure 7: Comparison of our PDS and SDGM [45] on CIFAR-10 generation. We report the FID score and the corresponding relative runtimes. Relative runtime is with respect to the vanilla SGM.
@@ 8 Conclusion
We have proposed a novel preconditioned diffusion sampling (PDS) method for accelerating off-the-shelf score-based generative models (SGMs), without model retraining. Considering the sampling process as a combination of Langevin dynamics and reverse diffusion, we reveal that existing sampling suffers from the ill-conditioned issue. To solve this, we reformulate both the Langevin dynamics and the reverse diffusion with statistics based matrix preconditioning. Theoretically, PDS preserves the steady-state (i.e., target) distribution of the Langevin dynamics and the final-state distribution of the reverse diffusion respectively. We show that PDS significantly accelerates existing state-of-the-art SGMs while maintaining the generation quality, matching or even surpassing representative SGMs and DDPMs.
Figure 9: Facial images generated by NCSN++ [4] with our PDS using different solenoidal items. Iterations $T$: $66$. Dataset: FFHQ [21] ($1024\times 1024$).
Figure 10: Facial images at a resolution of $1024\times 1024$ generated by NCSN++ [4] with our PDS. Sampling iterations: $66$. Dataset: FFHQ [21].
Appendix
Effect of PDS for alleviating ill-conditioning
We empirically investigate the effect of PDS to reduce the ill-conditioning of the original generation process of SGM. To measure the degree of ill-conditioning, we define the following to metrics: the coordinate variation and $V_{\mathrm{coo}}$ (variance across coordinates) and coordinate range $R_{\mathrm{coo}}$ (max coordinate value -min coordinate value). For the sample vector $\mathbf{x}\in\mathbb{R}^{d}$, the coordinate variation $V_{\mathrm{coo}}$ is define as
$\sum_{i=1}^{d}(\mathbf{x}_{i}-\frac{1}{d}\sum_{j=1}^{d}\mathbf{x}_{j})^{2},$
and the coordinate range $R_{\mathrm{coo}}$ is define as
$\max_{i}\mathbf{x}_{i}-\min\mathbf{x}_{i}.$
Both metrics indicate the gap of the scales in different coordinates, i.e., the degree of the ill-condition issue, with larger $R_{\mathrm{coo}}$ and $V_{\mathrm{coo}}$ indicate more severe ill-condition issue. We add new experiments, in which we generate some samples through NCSN++ with and without PDS on CIFAR-10, CelebA 64x64 and ImageNet and evaluate the above two metrics. Both metrics are average over all samples and iteration time points for each dataset. As shown in Fig. 11, PDS can reduce both coordinate variation and $V_{\mathrm{coo}}$ coordinate range $R_{\mathrm{coo}}$, hence alleviates the ill-condition issue.
Figure 11: The coordinate variation $V_{\mathrm{coo}}$ and coordinate range $R_{\mathrm{coo}}$ of NCSN++ [4] with PDS and without PDS for CIFAR-10, ImageNet and CelebA datasets. We set iterations $T=100$.
Details of implementation
@@
For CIFAR-10, LSUN (bedroom and church, $256\times 256$), and FFHQ ($1024\times 1024$), we apply the checkpoints of NCSN++ provided in [4]. We train NCSN++ on CelebA ($64\times 64$) for 600k iterations at the batch size of 64 on two NVIDIA RTX 3090 GPUs. We apply the same hyper-parameters of optimizer and learning schedule as [4]. For sampling, we use the same generation process of continuous variance exploding (VE)-NCSN++ [4] with both predictor and corrector, and all the experiments apply the linear schedule used in [4] with both predictor and corrector (PC).
Extreme low iteration generation
We investigate the performance of PDS under an extreme low iteration number. We find that the FID score of PDS is 304.71 for generating CIFAR-10 with 20 iterations. While this performance falls short of state-of-the-art VP SGMs such as DDIM, Analytic-DPM, and PNDM, it still surpasses the baseline NCSN++ model, which yields an FID score of 460.51. This suggests that our PDS still exerts a beneficial influence on the generation process, although it is unable to completely mitigate the issue of ill-conditioning at such low iteration numbers.
One plausible explanation for this phenomenon is the inferior numerical stability of VE SGMs, which includes both the vanilla NCSN++ and our PDS, compared to VP SGMs. Consequently, the PDS faces greater challenges when generating data with a limited number of iterations Nevertheless, it is essential to highlight that the both vanilla NCSN++ and our PDS can achieve superior performance at moderate or higher iteration number, compared to these VP SGMs. As shown in Tab. 2-4 in the main paper, our PDS achieves the best FID across a wide range of iterations, spanning from 100 to 1000. Consequently, our PDS presents a compelling improvement in the balance between generation quality and speed when compared to existing methods.
Figure 12: Sampling using NCSN [2] on MNIST. Top: Results by the original sampling method. Bottom: Results by our PDS ($(r,\lambda)=(0.2H,1.6)$). Iterations: $T=20$.
Figure 13: Sampling using NCSNv2 [3] on LSUN church ($96\times 96$). Left: Results by the original sampling method with different sampling iterations. Right: Results by our PDS ($(r,\lambda)=(0.2H,1.1)$) with different iterations $T$.
Figure 14: Sampling using NCSNv2 [3] on LSUN tower ($128\times 128$). Left: Results by the original sampling method with different sampling iterations. Right: Results by our PDS ($(r,\lambda)=(0.2H,1.6)$) with different iterations $T$.
More samples
We provide more samples generated by NCSN++ with PDS in Fig. 16-17.
Figure 15: Facial images generated by NCSN++ [4] with our PDS under different iterations $T$ and different solenoidal items described in Eq. (53)-(57). Dataset: FFHQ [21] ($1024\times 1024$).
@@
Figure 16: CIFAR-10 samples generated by NCSN++ [4] with PDS under different iterations $T$.
@@
Figure 17: CelebA ($64\times 64$) samples generated by NCSN++ [4] with PDS under different iterations $T$.

### v3

@@ 7.2 Evaluation on running speed
We evaluate the effect of running speed by using PDS. In this test, we use one NVIDIA RTX 3090 GPU. We track the average wall-clock time of generating a batch of 8 images. As shown in Table 8, PDS only introduces marginal extra time complexity per iteration, corroborating our complexity analysis in Sec. 5.3.3. Importantly, PDS enables low-iteration generation for better efficiency. For example, for the high-resolution generation task on FFHQ ($1024$), PDS reduces the time cost of generating $8$ human facial samples from $1920$ to $68$ in seconds, i.e., 28$\times$ speedup.
Table 8: Comparing the running time per iteration of generating a batch of $8$ images. Time unit: Seconds.
@@
7.3 Experiments on other SGMs
We evaluate the performance of PDS on two more SGMs, i.e., NCSN (Song and Ermon, 2019) and NCSNv2 (Song and Ermon, 2020). Although the approach of constructing preconditioning operators in Sec. 5.3 works well for accelerating NCSN++ (Song et al., 2021b), it is less effective for NCSN (Song and Ermon, 2019) and NCSNv2 (Song and Ermon, 2020). A possible reason is that the two models are less capable of exploiting the data statistics. To address this, we construct a simpler version of the frequency preconditioner as
$R_{f}(c,h,w)=\left\{\begin{matrix}1,\quad(h-H/2)^{2}+(w-W/2)^{2}\leq 2r^{2}\\ \lambda,\quad(h-H/2)^{2}+(w-W/2)^{2}>2r^{2}\end{matrix}\right.,$ (37)
$M_{f}[\mathbf{x}]=F^{-1}[F[\mathbf{x}]\bullet R_{f}].$ (38)
Note that NCSN and NCSNv2 have only correctors in the sampling process. We set $\lambda$ so that the operator $M_{f}$ balances the coordinates of low and high-frequency components of its input via shrinking the scale of low-frequency coordinates. This is because that the scale of low-frequency coordinates is higher than that of high-frequency coordinates among natural images (van der Schaaf and van Hateren, 1996). As shown in Fig. 6-8, PDS can dramatically accelerate the sampling process of NCSN and NCSNv2.
Figure 6: Sampling using NCSN (Song and Ermon, 2019) on MNIST. Left: Results by the original sampling method. Right: Results by our PDS ($(r,\lambda)=(0.2H,1.6)$). Iterations: $T=20$.
Figure 7: Sampling using NCSNv2 (Song and Ermon, 2020) on LSUN church ($96\times 96$). Left: Results by the original sampling method with different sampling iterations. Right: Results by our PDS ($(r,\lambda)=(0.2H,1.1)$) with different iterations $T$.
Figure 8: Sampling using NCSNv2 (Song and Ermon, 2020) on LSUN tower ($128\times 128$). Left: Results by the original sampling method with different sampling iterations. Right: Results by our PDS ($(r,\lambda)=(0.2H,1.6)$) with different iterations $T$.
7.4 Compared with subspace diffusion generative models
The subspace diffusion generative models (SDGMs) (Jing et al., 2022) restrict the sampling process of the SGMs on a low-dimensional subspace for acceleration. This might help mitigate the issue of ill-conditioning. We thus compare our PDS with SDGMs in terms of the effect of alleviating ill-conditioning.
We conduct two separate experiments to evaluate their performance. In the first experiment, we investigate the performance of two methods in terms of varying iterations. As shown in Table 9, our PDS achieves lower/better FID scores when compared to SDGMs subject to the same iterations. It is important to highlight that SDGM benefits from a smaller per-iteration cost due to reduced dimensionality, allowing it to execute more iterations within the same runtime. Hence we conduct a second experiment with a comparison taking into account the runtime and sample quality of the two methods, a different fair evaluation. As depicted in Fig. 9, our PDS demonstrates the ability to generate higher-quality images (as indicated by a lower FID score) at the equivalent runtime cost when compared to SDGMs. In summary, our preconditioning strategy surpasses the dimension reduction idea of SDGM.
Table 9: Comparison of our PDS and SDGM (Jing et al., 2022) on CIFAR-10 generation under varying iterations.
Iterations $T$ 50 100 200 400 500 1000
PDS 4.90 2.90 2.53 2.19 2.11 1.99
SDGM 386.70 64.05 7.03 3.03 2.70 2.17
Figure 9: Comparison of our PDS and subspace diffusion generative models (SDGM) (Jing et al., 2022) on CIFAR-10. We report the FID score and the relative runtimes with respect to the vanilla SGM.
8 Further experiments
8.1 Effect of PDS for alleviating ill-conditioning
We empirically investigate the effect of PDS to reduce the ill-conditioning of the original generation process of SGM. To measure the degree of ill-conditioning, we define the following metrics: the coordinate variation and $V_{\mathrm{coo}}$ (variance across coordinates) and coordinate range $R_{\mathrm{coo}}$ (max coordinate value -min coordinate value). For the sample vector $\mathbf{x}\in\mathbb{R}^{d}$, the coordinate variation $V_{\mathrm{coo}}$ is define as $V_{\mathrm{coo}}=\sum_{i=1}^{d}(\mathbf{x}_{i}-\frac{1}{d}\sum_{j=1}^{d}\mathbf{x}_{j})^{2}$, and the coordinate range $R_{\mathrm{coo}}$ is define as $R_{\mathrm{coo}}=\max_{i}\mathbf{x}_{i}-\min\mathbf{x}_{i}$. Both metrics indicate the gap of the scales in different coordinates, i.e., the degree of the ill-condition issue, with larger $R_{\mathrm{coo}}$ and $V_{\mathrm{coo}}$ indicating more severe ill-condition issue. We add new experiments, in which we generate some samples through NCSN++ with and without PDS on CIFAR-10, CelebA 64x64, and ImageNet and evaluate the above two metrics. Both metrics are average over all samples and iteration time points for each dataset. As shown in Fig. 10, PDS can reduce both coordinate variation and $V_{\mathrm{coo}}$ coordinate range $R_{\mathrm{coo}}$, hence alleviating the ill-condition issue.
Figure 10: The coordinate variation $V_{\mathrm{coo}}$ and coordinate range $R_{\mathrm{coo}}$ of NCSN++ (Song et al., 2021b) with PDS and without PDS for CIFAR-10, ImageNet and CelebA with iterations $T=100$.
@@ 8.2 Relation between $\alpha$ and $T$
We test CIFAR-10 with only the frequency preconditioner and CelebA, ImageNet, and COCO ($64\times 64$) with both. The results of linear regression are depicted in Fig. 11. We obtain the coefficient of determination of 0.895 on CIFAR-10, 0.940 on CelebA($64\times 64$), 0.992 on ImageNet ($64\times 64$), and 0.804 respectively. This suggests strong linear relations, validating our analysis. Note, when $(1+\frac{1}{\alpha})^{2}-1$ and $(1+\frac{1}{\alpha})-1$ are similar, the ratio of the slope $a$ will be close to the square root of the dimension ratio. For instance, the relation between CIFAR-10 and CelebA($64\times 64$) fits as
$\frac{a_{1}}{a_{2}}=\frac{28.53}{50.67}=0.56\sim 0.5=\sqrt{\frac{3\times 32\times 32}{3\times 64\times 64}}=\sqrt{\frac{d_{1}}{d_{2}}},$
@@
Figure 11: Relation between the iterations $T$ and the parameter $\alpha$. All cases suggest a strong linear correlation.
@@ 8.3 Sensitivity of parameters
We evaluate the parameter sensitivity of PDS. We perturb the optimal $\alpha$ (Table 1) by ±$\pm 10\%$. As shown in Table 10, the performance is not parameter-sensitive.
Table 10: Sensitivity analysis on of the parameter($\alpha$) on the CIFAR-10.
@@
8.4 Skew-symmetric operator
——–$\omega=1$———–$\omega=10$———-$\omega=100$———-$\omega=1000$
Figure 12: Samples produced by NCSN++ (Song et al., 2021b) w/ PDS on FFHQ ($1024$x1024102410241024(Karras et al., 2019) with different solenoidal terms. Iterations: $66$.
We study the effect of the solenoidal term $S\bigtriangledown_{\mathbf{x}}\log p_{\mathbf{x}^{\ast}}(\mathbf{x})$ to the Langevin dynamics. As proved in Thm. 4.1, as long as $S$ is skew-symmetric, it will not change the steady-state distribution of the original process. To verify this, we consider the process
@@ 8.4 Skew-symmetric operator
where $\omega$ controls the weight of $S$. We experiment with $S[\cdot]=\mathrm{Re}[F[\cdot]-F^{\mathrm{T}}[\cdot]]$ for being skew-symmetric and $\omega\in[1,1000]$. Fig. 12 shows that $\omega$ has no impact on the generation quality. This verifies that $S$ does not change the original steady-state distribution. We further perform similar tests with different iterations and other different skew-symmetric operator
@@
where $P_{m,n}$ is the shift operator that rolls the input image for $m$ places along the height coordinate and rolls the input image for $n$ places along the width coordinate. The sampling results are shown in Fig. 13, where we set $\omega=1000$. It is observed that again all these solenoidal terms do not impose an obvious effect on the sampling quality. Additionally, as displayed in Fig. 5 in the supplementary material, these solenoidal terms also do not have an obvious effect on acceleration. We leave a more extensive study of the solenoidal terms in the further work.
Figure 13: Facial images generated by NCSN++ (Song et al., 2021b) with our PDS using different solenoidal items. Iterations $T$: $66$. Dataset: FFHQ ($1024\times 1024$(Karras et al., 2019).
@@ 9 Conclusion
We have proposed a novel preconditioned diffusion sampling (PDS) method for accelerating off-the-shelf score-based generative models (SGMs), without model retraining. Considering the sampling process as a combination of Langevin dynamics and reverse diffusion, we reveal that existing sampling suffers from the ill-conditioned issue. To solve this, we reformulate both the Langevin dynamics and the reverse diffusion with statistics-based matrix preconditioning. Theoretically, PDS preserves the steady-state (i.e., target) distribution of the Langevin dynamics and the final-state distribution of the reverse diffusion respectively. We show that PDS significantly accelerates existing state-of-the-art SGMs while maintaining the generation quality, matching or even surpassing representative SGMs and DDPMs.
Appendix A Implementations
@@
For CIFAR-10, LSUN (bedroom and church, $256\times 256$), and FFHQ ($1024\times 1024$), we apply checkpoints of NCSN++ provided in (Song et al., 2021b). We train NCSN++ on CelebA ($64\times 64$) for 600k iterations at the batch size of 64 on two NVIDIA RTX 3090 GPUs. We apply the same hyper-parameters of optimizer and learning schedule as (Song et al., 2021b). For ImageNet and COCO datasets, we train the NCSN++ (Song et al., 2021b) by ourselves at the resolution of 64$\times$64 as the baseline. We apply the same hyper-parameters of optimizer and learning schedule as (Song et al., 2021b). We train the model for 2400k iterations for ImageNet and 600k for COCO. The batch size is set at 256 for both datasets. For sampling, we use the same generation process of continuous variance exploding (VE)-NCSN++ (Song et al., 2021b) with both predictor and corrector and all the experiments apply the linear schedule used in (Song et al., 2021b) with both predictor and corrector (PC).
Appendix B Experiments on latent space
We assess the performance of our PDS on SGMs in latent spaces. Since existing latent SGMs predominantly rely on VP SDE or ODE frameworks for generation, which are incompatible with PDS, we train custom latent SGMs. For the MNIST dataset, we first train a convolutional variational autoencoder (VAE) to map images into a 384-dimensional latent space, then train an NCSN on the VAE-transformed dataset. Similarly, for the CelebA ($64\times 64$) dataset, we use a convolutional VAE to transform images into a 512-dimensional latent space and train an NCSN on this transformed dataset. We apply PDS with $\alpha=1.75$ for MNIST and $\alpha=4$ for CelebA. As shown in Fig. 14 and Fig. 15, PDS substantially preserves sampling quality, while achieving 20×20$\times$20 × and 10×10$\times$10 × acceleration for the two datasets, respectively.
The latent SGM serves as a nonlinear counterpart to the proposed PDS. While latent SGMs rely on a nonlinear VAE for sampling space transformation, PDS employs a linear operator to achieve similar effects. Both methods address ill-conditioning, but they differ in execution: latent SGMs utilize the dimensionality reduction of a VAE, whereas PDS directly applies an efficient linear operator. Latent SGMs, however, require training a VAE and subsequently retraining the SGM for acceleration, whereas PDS bypasses this step by leveraging dataset statistics, enabling seamless integration with existing SGMs without retraining or additional VAE training. Furthermore, experimental results (Fig. 14 and Fig. 15) indicate that PDS enhances the sampling quality of latent SGMs. This suggests that the VAE does not fully resolve the ill-conditioning inherent in the latent SGM sampling process.
@@ Appendix B Experiments on latent space
Figure 14: Images generated by NCSN (Song et al., 2021b) on the latent MNIST with or without our PDS under different iterations $T$.
@@
Figure 15: Images generated by NCSN (Song et al., 2021b) on the latent CelebA ($64$) with or without our PDS under different iterations $T$.
Appendix C More samples
We provide more samples generated by NCSN++ with PDS in Fig. 16-22.